In a steam nozzle steam expands from 15 bar to 5 bar with initial temperature of 300°C and mass flow of 1 kg/s. Determine the throat and exit areas considering (i) expansion to be frictionless and, (ii) friction loss of 8% throughout the nozzle. 


1
Expert's answer
2021-06-10T03:46:53-0400

Initial pressure(p1) = 15 bar

Initial Temperature(T1)= 300°C

Exit pressure(p2) = 5 bar

Nozzle efficiency (nozzle) = 100%


mass flow of steam (v1)= 1kg/s

Discharge mass flow (v2) = ?


From steam table, For 15 bar, 300°C, h1 = 2943 kJ/kg s1 = 6.926 kJ/kg.K

From steam table, For 5 bar,

hf2 = 439.3 kJ/kg ; hfg2 = 2244.1 kJ/kg;

sf2 = 1.3 61 kJ/kg K ; sfg2 = 5.937 kJ/kgK.


Since "s_1=s_2,"

s1 = sf2 + x2sfg2

6.926 = 1.361 + x2(5.937)

x2 =0 .9373


We know that,

h2 = hf2+x2hfg2

h2 = 439.3+(0.9373)2244.1

h2 = 2542KJ/Kg


Exit velocity (v2) = Rt[(2000(2943) 2542) + 50²] = 896.91kg/s.